the following costs were incurred by odd a mfg cost

DM =105200

DL = 221400

MOH = 98600

based on the above data

a.determine prime cost

b.conversion cost

c.mfg cost


1
Expert's answer
2021-08-18T15:54:57-0400

a. Prime Cost = Direct Material + Direct Labor

Prime cost = 105200 + 221400

Prime Cost = 326600


b. Conversion Cost = Direct Labor + Manufacturing overhead

Conversion Cost = 221400 + 98600

Conversion Cost = 320000


c. Manufacturing Cost = Direct Material + Direct Labor + Manufacturing Overhead

Manufacturing Cost = 105200 + 221400 + 98600

Manufacturing Cost = 425200